OpenStudy (anonymous):

S=π (1.5) √((1.5)^2+(4)^2) how to you solve this to get the answer?

OpenStudy (anonymous):

S=π 1.5 √(1.5^2+4^2) actually its like this

OpenStudy (anonymous):

just break it into steps: \[1.5^2 + 4^2\] then sqrt it. then multiply it by 1.5 and then if you are looking for a decimal answer multiply in your \[\pi\]

OpenStudy (anonymous):

I tried that but I keep getting the wrong answer? one time I got 4.6 then another time 0 and the answer is supposed to be 20.13n^2

OpenStudy (anonymous):

Okay, what do you get when you plug in \[1.5^2\] and \[4^2\] and when you add them.

OpenStudy (anonymous):

18.25

OpenStudy (anonymous):

yeah, now whats the sqrt of that?

OpenStudy (anonymous):

333.06

OpenStudy (anonymous):

Err I meant the square root of 18.25\[\sqrt(18.25)\] not \[18.25^2\]

OpenStudy (anonymous):

4.3

OpenStudy (anonymous):

right, now take that times 1.5 and \[\pi\]

OpenStudy (anonymous):

ok how would I write that Im not doing something write I came up with zero again

OpenStudy (anonymous):

plug \[4.3* \pi *1.5 \]into your calculator

OpenStudy (anonymous):

got 6.45\[\pi\]

OpenStudy (anonymous):

which is correct, you could also use 3.14 * \[\pi\]

OpenStudy (anonymous):

Ok Im confused then how do I get the 20.13n^2 answer